OpenStudy (anonymous):

Convert the equation 12x – 3y = 6 to slope-intercept form.

OpenStudy (anonymous):

A. y = 4x + 2 B. y = 4x – 2 C. y = 12x – 6 D. y = -4x – 2

OpenStudy (anonymous):

refer to this link http://openstudy.com/study?login#/updates/4fa14748e4b029e9dc321fe8

OpenStudy (anonymous):

\[y=mx+b\] Solve for y. m is your slope, b is your y-intercept.

OpenStudy (anonymous):

I believe the answer is y = 4x - 2